# Reliably Detecting Model Failures in Deployment Without Labels

### v3

@@ 1 Introduction
Performance guarantees of conventional machine learning (ML) models hinge on the belief that the distribution of data with which these models train is identical to the distribution on which they are deployed (Rabanser et al., 2019; Recht et al., 2018; Santurkar et al., 2020). In many real-world scenarios such as healthcare, however, this assumption fails due to distribution shift during model deployment. Benchmarks such as WILDS (Koh et al., 2021) and WILD-Time (Yao et al., 2022) have encouraged machine learning researchers to study and better understand how data shifts influence predictive systems. Yet, the number of tools at a practitioner’s disposal for building predictive models far exceeds those to monitor model failures. There is a need to create guardrails that self-detect and alert end-users to critical changes in the model when its performance drops below acceptable thresholds (Habib et al., 2021; Zadorozhny et al., 2022).
@@
We define post-deployment deterioration (PDD) as the scenario where a trained ML model underperforms on a distributionally shifted deployment query with respect to its validation performance. PDD presents a distinct set of systemic challenges stemming from considerations over the feasibility of deployment in real-world ML pipelines. Predominant is the scarcity of labels during deployment: for many downstream tasks such as in healthcare, labels are expensive to obtain (Razavian et al., 2015) or require human intervention (Liu et al., 2022). Due to deployed models predicting events temporally extended in the future (Boodhun and Jayabalan, 2018; Zhang et al., 2019), labels might even be unavailable. Another is the robustness of the monitoring system: it should flag critical changes in model deterioration early, using few samples, while remaining resilient to non-deteriorating changes to minimize unnecessary interruptions of service among other practical considerations.
@@
To address these challenges, we conceive a set of desiderata for any algorithm monitoring PDD, targeting their practicality and effectiveness as plug-ins to ML pipelines. To address the scarcity of labels, PDD monitoring algorithms should operate on unlabeled data from the test distribution to ascertain potential deterioration of the deployed model. Further, PDD monitoring algorithms should not depend on training data during deployment, as continuous (even indefinite) access to sensitive or personally identifiable training data might violate certain regulations protecting the privacy of data subjects (Mühlhoff, 2023). An algorithm satisfying this desideratum is scalable, as it only audits a model’s input stream during monitoring with minimum data storage and regulatory concerns. Finally, PDD monitoring mechanisms should be robust to flagging non-deteriorating changes and effective even when samples from the deployment distribution are scarce.
@@
When it comes to designing monitoring protocols that satisfy the above desiderata, recent related works only partially attend to individual desiderata. The literature on distribution shifts while achieving strong empirical performance on unlabeled deployment data (Liu et al., 2020; Zhao et al., 2022; Zhang et al., 2024), are not robust to false positives when the distribution shift is non-deteriorating. The model disagreement framework (Yu and Aizawa, 2019; Chuang et al., 2020; Jiang et al., 2021; Ginsberg et al., 2023; Rosenfeld and Garg, 2023) emerges as a competitive setup for monitoring with downstream performance considerations via the tracking of disagreement statistics, while foregoing explicit distribution shift computations. However, shift-based and disagreement-based monitoring methods all depend on the presence of training data post-deployment, and do not provide any guarantees on robustness against false positives in the monitoring of non-deteriorating shifts.
@@
Empirical validation. We showcase experimental results on various shift scenarios in the UCI Heart Disease dataset (Janosi et al., 1989), CIFAR-10/10.1 (Recht et al., 2018), Camelyon17 (WILDS) (Koh et al., 2021), and the GEneral Medicine INpatient Initiative (GEMINI) dataset  (Verma et al., 2017, 2021) to demonstrate its effectiveness in monitoring models of various modalities. Our method effectively detects deployment-time deterioration with low false positive rates (FPR) when shifts are non deteriorating, and achieves competitive true positive rates (TPR) when shifts are deteriorating compared to standard baselines. In particular, we discuss how results on the internal medicine dataset suggest D3M to be well-suited for integration into real-world clinical monitoring pipelines.
@@
  Training data-free Deteriorating Non-deteriorating Disagreement         Yu and Aizawa, 2019 Yu and Aizawa (2019) ✓         Liu et. al., 2020 Liu et al. (2020) ✗         Jiang et. al., 2021 Jiang et al. (2021) ✓         Zhao et. al., 2022 Zhao et al. (2022) ✗         Rosenfeld and Garg, 2023 Rosenfeld and Garg (2023) ✓         Ginsberg et. al., 2023 Ginsberg et al. (2023) ✓         D3M (ours) ✓
@@ 2.1 Problem setup
Leveraging insights from Yu and Aizawa (2019); Goldwasser et al. (2020); Ginsberg et al. (2023), the framework of model disagreement possesses this property under certain assumptions about the underlying distribution change (see Appendix A). We say that two models $h_{1}$ and $h_{2}$ disagree on an input $x\in\mathcal{X}$ if $h_{1}(x)\not=h_{2}(x)$. In particular, models exhibit greater predictive disagreement on unsupervised samples that lead to model deterioration, compared to in-distribution (ID) samples. This is observed through the increased entropy-based discrepancy between classification heads in Yu and Aizawa (2019), or maximum disagreement between models in the same ensemble in Goldwasser et al. (2020) and Ginsberg et al. (2023), as signal for detecting deployment deterioration. Maximizing classification head discrepancy for OOD detection (Yu and Aizawa, 2019) is efficient for monitoring at deployment time, requiring only one forward pass to compute a verdict. However, this trades off classification accuracy as this training procedure alters the original trained decision boundaries. On the other hand, computing model disagreement between ensembles (Ginsberg et al., 2023) requires finetuning a potentially large network to collect ID and deployment-time disagreement statistics $\phi$. In addition, ID training data is required at deployment, further limiting the scalability of such framework.
@@ 2.2 Overview of D3M
1. (Train) Base model training. Firstly, a neural feature extractor $\operatorname{FE}_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$ coupled with a Variational Bayesian Last Layers (Harrison et al., 2024) $\operatorname{VBLL}_{\theta}:\mathbb{R}^{d}\to\mathcal{P}(\mathbb{R}^{C})$ parametrized by $\theta\in\Theta$, are trained on supervised ID data $\mathcal{D}^{n}\sim\bm{P}^{n}$ to output posteriors over logits corresponding to $C$ classes. For an input $x\in\mathcal{X}$ and its ground-truth label $y\in\mathcal{Y}$:
@@ 2.3 Algorithmic insights
D3M is free from training data. D3M is designed to monitor deployment-time input streams without requiring the training data of its base model. This is an important advantage shared by Yu and Aizawa (2019), however not enjoyed by other baselines in Table 1, whose computation of $\phi$ statistics require maintaining agreement on training data. When the neural feature extractor $\operatorname{FE}$ is millions or billions of parameters, trained with a comparably large dataset, it becomes infeasible to store the training data within edge compute nodes in order to run the monitoring mechanism.
No tradeoff with prediction accuracy. Furthermore, D3M does not trade-off prediction accuracy as the mean prediction model from the PPD is not modified during the construction of $\Phi$ nor its deployment (Appendix C.1). Therefore, ID generalization theories remain applicable (Vapnik, 1999; Bartlett and Mendelson, 2002; Zhang et al., 2016; Nagarajan and Kolter, 2019; Jiang et al., 2019; Lee et al., 2020). In addition, recording a collection of ID maximum disagreement rates reinforces the robustness of the method. While single instances of maximum disagreement rate are subject to noise in the choice of the held-out validation set, the sampling of posterior logits $z_{i}^{(k)}$, and the number of drawn samples $K$, a collection $\Phi$ of maximum disagreement rates coupled with a quantile test softens the effect of noise by leveraging large sample statistics.
Less diverse samples compared to its fully Bayesian counterpart. Harrison et. al., 2024 Harrison et al. (2024) suggests that one may apply a variational Bayesian treatment to the last layer only, saving on computational costs of forward and backward propagating through a Bayesian neural network, and enjoy largely the same uncertainty estimation and OOD detection performance. However, the usage of VBLL in D3M is unorthodox with respect to the original work: by sampling posterior weights from a Bayesian model, we are hoping to land on a set of weights such that the resulting model strongly disagree with the base model. VBLL allows more efficient sampling compared to to a fully Bayesian network, though lacking diversity in the sampled predictions, as the variability comes from the last layer only. This results in $K$ sampled logits mostly agreeing with the mean prediction. To improve the diversity of sampling, we employ (1) temperature scaling and (2) the sampling of predictions from the categorical distribution for the computation of maximum disagreement rather than argmax labeling. We defer the study of this tradeoff between fully Bayesian and VBLL to Appendix C.5.
@@ 2.4 Implementing D3M
Two core limit The composition $\operatorname{VBLL}_{\theta}\circ\operatorname{FE}_{\theta}:\mathcal{X}\to\mathcal{P}(\mathbb{R}^{C})$ is readily trained end-to-end on ID data of size $n$ by maximizing the ELBO. For rounds $t\in[T]$, on a large held-out ID validation set, we sample datasets $\mathcal{D}^{m}_{t}$ of size $m\ll n$ with replacement for calibration. $\mathcal{D}^{m}_{t}$ is processed in a single forward pass, producing $m$ independent $C$-dimensional Gaussian posteriors:
@@ 3.1 Experimental setup
Datasets. (1) The UCI Heart Disease prediction dataset (Janosi et al., 1989), where each hospital corresponds to a different domain. Here, the distribution shift is due to differences in patient populations and data collection practices across hospitals. (2) CIFAR-10/10.1 datasets (Krizhevsky et al., 2009; Recht et al., 2019) where shift comes from subtle changes in the dataset creation process. By viewing samples from CIFAR-10 as $\bm{P}$, we test our models’ ability to flag deteriorating shift from samples in $\bm{Q}=$ CIFAR-10.1. (3) The Camelyon17 dataset from the WILDS benchmark (Koh et al., 2021; Yao et al., 2022) a histopathology image dataset for detecting metastases in lymph node slides, where distribution shift arises from variations in slide staining and image acquisition between hospitals contributing the data. (4) The General Medicine Inpatient Initiative (GEMINI) (Verma et al., 2017, 2021) dataset, a comprehensive repository of standardized clinical and administrative data from hospitalizations within general internal medicine. We focus on predicting patient mortality within a 14-day horizon, leveraging static and longitudinal clinical features. This task is deemed essential for facilitating timely clinical interventions, optimizing the allocation of healthcare resources, and ultimately striving to improve patient outcomes Naemi et al. (2021); Barsasella et al. (2022); Hernández Guillamet et al. (2023). Detailed data descriptions can be found in B.3. Full sweeping details and hyperparameter configurations are reported in B.5.
Implementation & Baselines. For tabular data (UCI, GEMINI 14-day mortality), in our implementation, $\operatorname{FE}_{\theta}$ corresponds to sequences of affine transformations and nonlinearities with skip-connections. For image datasets (CIFAR-10/10.1, Camelyon17), $\operatorname{FE}_{\theta}$ is either a trained or a finetuned ResNet (He et al., 2016). In particular, the D3M mechanism is agnostic to the choice of feature extractor, provided that the architecture reflects appropriate inductive biases for the data modality and permits gradient flow through the extracted features.
To demonstrate that our D3M enjoys low FPR on non deteriorating shifts and high TPR on deteriorating shifts, we compare it against several distribution divergence-based detection methods from the literature: Deep Kernel MMD (MMD-D) (Liu et al., 2020), H-divergence (Zhao et al., 2022), Black Box Shift Detection (BBSD) (Lipton et al., 2018), Relative Mahalanobis Distance (RMD) Ren et al. (2021), Deep Ensembles (Ovadia et al., 2019), CTST (Lopez-Paz and Oquab, 2016), Domain Classifier (DC) (Rabanser et al., 2019), and Detectron Ginsberg et al. (2023). Details can be found in Appendix B.1.
@@ 3.2 Discussions and Analysis
Comparisons with Detectron. Ginsberg et al. (2023) While D3M and Detectron achieve similar monitoring performance—both successfully flagging all deteriorating shifts in higher query size scenarios (see B.6)—D3M is significantly more scalable and practical for real-world deployment, especially at the edge (Appendix C.2). Unlike Detectron, which requires persistent access to the original training data and gradient-based finetuning during deployment, D3M operates in a truly “source-free” fashion post-training: it needs neither storage nor access to sensitive training samples, nor does it perform computationally expensive or potentially destabilizing finetuning in production. Furthermore, by relying solely on forward passes and simple statistics over deployment data, D3M is inherently robust and less susceptible to the risks associated with continual retraining, such as accidental data contamination or privacy leakage.
@@ 4 Related Work
Adaptation. Test-time adaptation (TTA) concerns itself more with how one achieves strong deployment performance in spite of a critical change, rather than how one would flag this critical change. To this end, we fundamentally believe that monitoring offers an additional level of flexibility: when a deterioration flag is raised, adaptation is a possible course of action among many others. A few works from this rich body of literature are tangentially related to disagreement-based monitoring. Saito et al. (2018); Vu et al. (2019); Wang et al. (2020) all leverage prediction uncertainty or discrepancy to drive unsupervised domain adaptation. Zhang and Chen (2023), measures and minimizes prediction disagreement at test time, adaptively aligning target-domain features to a domain-invariant feature space, improving performance without requiring access to target domain data during training. In the continual learning literature, Ye and Bors (2022) also leverages classifier disagreement as a unsupervised proxy of distributional change.
@@
Performance monitoring of ML models & deteriorating shift. Evaluating a model’s reliability during deployment is crucial for the safety and effectiveness of the machine learning pipeline over time. Feng et al. (2024b, a) provided a causal viewpoint wherein the challenge to adapt to diverse scenarios still remain due to a lack of access to the true causal graph. Model disagreement is often used as a monitoring tool for the model generalization (Jiang et al., 2022; Chuang et al., 2020; Ginsberg et al., 2023; Rosenfeld and Garg, 2023). These works align strongly with our method, though the framing is orthogonal to ours as our analysis provides FPR and TPR guarantees whereas they provided sufficient conditions in either ID or other shifts beyond our scope. Several works in the recent literature differentiate shifts in terms of deteriorating or non-deteriorating shifts. Podkopaev and Ramdas (2021) studied (deteriorating) shift detection in the continuous monitoring setting using a sequential hypothesis test. Due to the setting being sequential in nature, their method requires true labels from $\bm{Q}$ immediately after prediction or at the least in a delayed fashion. Ginart et al. (2022) approached the monitoring problem from a time-continuous anomaly detection perspective, allowing periodic querying of deployment-time ground truths from experts. Other related empirical works along this literature are Wang et al. (2023); Kamulete (2022), and the very applied Nigenda et al. (2022).
@@
Distribution shift detection. Methods to detect distribution shift arise from different perspectives. In covariate shift detection, (Lopez-Paz and Oquab, 2016; Liu et al., 2020; Zhao et al., 2022) treated detection as two-sample tests via classifier, Deep Kernel MMD, and H-divergence. For label shift on the other hand, (Lipton et al., 2018; Azizzadenesheli et al., 2019) formulated the problem as a convex optimization problem by solving the label distribution ratio $\alpha=\bm{Q}(y)/\bm{P}(y)$. The problem of OOD detection (Liang et al., 2017; Kamulete, 2022) seeks to detect if an individual sample $x$ comes from the training distribution $x\sim\bm{P}(x)$. Some previous works (Ren et al., 2019; Morningstar et al., 2021) also adopted the methods in covariate shift detection and generalization by estimating the density ratio for the identification of OOD samples. Whilst these methods detect shifts, they are constrained by their requirement of training data post-deployment and do not consider the extent to which shifts affect model performance.
@@
Estimating test error with unlabeled data. Another rich body of research is the estimation of (OOD) test error. This technique and its variants are often inspired by domain adaptation theories (Ben-David et al., 2006, 2010; Acuna et al., 2021; Ganin et al., 2016), seeking guarantess in the form of $\operatorname{{err}}(f;\bm{Q}_{g})\leq\operatorname{{err}}(f;\bm{P}_{g})+\Delta(f,\mathcal{H})$, with $\Delta(f,\mathcal{H})=\sup_{h\in\mathcal{H}}|\operatorname{{err}}(h;\bm{P}_{f})-\operatorname{{err}}(h;\bm{Q}_{f})|$. This objective can be alternatively viewed as searching for a critic function $h\in\mathcal{H}$ to maximize a performance gap (Rosenfeld and Garg, 2023; Jiang et al., 2021). One could thus provably estimate the upper bound of the test distribution error. These theories also implicitly assume the availability of training data. Further, they assume that test error should be larger than the training error (granted this is often the case for deteriorating OOD), making them sensitive to non deteriorating shifts as well i.e., a high FPR in detection. Our theoretical analysis addresses this gap.

### v4

@@ 1 Introduction
Performance guarantees of conventional machine learning (ML) models hinge on the belief that the distribution of data with which these models train is identical to the distribution on which they are deployed (rabanser2019failing, ; recht2018cifar, ; santurkar2020breeds, ). In many real-world scenarios such as healthcare, however, this assumption fails due to distribution shift during model deployment. Benchmarks such as WILDS (koh2021wilds, ) and WILD-Time (yao2022wild, ) have encouraged machine learning researchers to study and better understand how data shifts influence predictive systems. Yet, the number of tools at a practitioner’s disposal for building predictive models far exceeds those to monitor model failures. There is a need to create guardrails that self-detect and alert end-users to critical changes in the model when its performance drops below acceptable thresholds (habib2021epic, ; zadorozhny2022out, ).
@@
We define post-deployment deterioration (PDD) as the scenario where a trained ML model underperforms on a distributionally shifted deployment query with respect to its validation performance. PDD presents a distinct set of systemic challenges stemming from considerations over the feasibility of deployment in real-world ML pipelines. Predominant is the scarcity of labels during deployment: for many downstream tasks such as in healthcare, labels are expensive to obtain (razavian2015population, ) or require human intervention (liu2022deep, ). Due to deployed models predicting events temporally extended in the future (boodhun2018risk, ; zhang2019time, ), labels might even be unavailable. Another is the robustness of the monitoring system: it should flag critical changes in model deterioration early, using few samples, while remaining resilient to non-deteriorating changes to minimize unnecessary interruptions of service among other practical considerations.
@@
To address these challenges, we conceive a set of desiderata for any algorithm monitoring PDD, targeting their practicality and effectiveness as plug-ins to ML pipelines. To address the scarcity of labels, PDD monitoring algorithms should operate on unlabeled data from the test distribution to ascertain potential deterioration of the deployed model. Further, PDD monitoring algorithms should not depend on training data during deployment, as continuous (even indefinite) access to sensitive or personally identifiable training data might violate certain regulations protecting the privacy of data subjects (muhlhoff2023predictive, ). An algorithm satisfying this desideratum is scalable, as it only audits a model’s input stream during monitoring with minimum data storage and regulatory concerns. Finally, PDD monitoring mechanisms should be robust to flagging non-deteriorating changes and effective even when samples from the deployment distribution are scarce.
@@
When it comes to designing monitoring protocols that satisfy the above desiderata, recent related works only partially attend to individual desiderata. The literature on distribution shifts while achieving strong empirical performance on unlabeled deployment data (liu2020learning, ; zhao2022comparing, ; zhang2024test, ), are not robust to false positives when the distribution shift is non-deteriorating. The model disagreement framework (yu2019unsupervised, ; chuang2020estimating, ; jiang2021assessing, ; ginsberg2023a, ; rosenfeld2024almost, ) emerges as a competitive setup for monitoring with downstream performance considerations via the tracking of disagreement statistics, while foregoing explicit distribution shift computations. However, shift-based and disagreement-based monitoring methods all depend on the presence of training data post-deployment, and do not provide any guarantees on robustness against false positives in the monitoring of non-deteriorating shifts.
@@
Empirical validation. We showcase experimental results on various shift scenarios in the UCI Heart Disease dataset (heart_disease_45, ), CIFAR-10/10.1 (recht2018cifar, ), Camelyon17 (WILDS) (koh2021wilds, ), and the GEneral Medicine INpatient Initiative (GEMINI) dataset  (verma2017, ; verma2021assessing, ) to demonstrate its effectiveness in monitoring models of various modalities. Our method effectively detects deployment-time deterioration with low false positive rates (FPR) when shifts are non deteriorating, and achieves competitive true positive rates (TPR) when shifts are deteriorating compared to standard baselines. In particular, we discuss how results on the internal medicine dataset suggest D3M to be well-suited for integration into real-world clinical monitoring pipelines.
@@
  Training data-free Deteriorating Non-deteriorating Disagreement         Yu and Aizawa, 2019 yu2019unsupervised ✓         Liu et. al., 2020 liu2020learning ✗         Jiang et. al., 2021 jiang2021assessing ✓         Zhao et. al., 2022 zhao2022comparing ✗         Rosenfeld and Garg, 2023 rosenfeld2024almost ✓         Ginsberg et. al., 2023 ginsberg2023a ✓         D3M (ours) ✓
@@ 2.1 Problem setup
Leveraging insights from yu2019unsupervised ; goldwasser2020beyond ; ginsberg2023a , the framework of model disagreement possesses this property under certain assumptions about the underlying distribution change (see Appendix A). We say that two models $h_{1}$ and $h_{2}$ disagree on an input $x\in\mathcal{X}$ if $h_{1}(x)\not=h_{2}(x)$. In particular, models exhibit greater predictive disagreement on unsupervised samples that lead to model deterioration, compared to in-distribution (ID) samples. This is observed through the increased entropy-based discrepancy between classification heads in yu2019unsupervised , or maximum disagreement between models in the same ensemble in goldwasser2020beyond and ginsberg2023a , as signal for detecting deployment deterioration. Maximizing classification head discrepancy for OOD detection (yu2019unsupervised, ) is efficient for monitoring at deployment time, requiring only one forward pass to compute a verdict. However, this trades off classification accuracy as this training procedure alters the original trained decision boundaries. On the other hand, computing model disagreement between ensembles (ginsberg2023a, ) requires finetuning a potentially large network to collect ID and deployment-time disagreement statistics $\phi$. In addition, ID training data is required at deployment, further limiting the scalability of such framework.
@@ 2.2 Overview of D3M
1. (Train) Base model training. Firstly, a neural feature extractor $\operatorname{FE}_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$ coupled with a Variational Bayesian Last Layers (harrison2024variational, ) $\operatorname{VBLL}_{\theta}:\mathbb{R}^{d}\to\mathcal{P}(\mathbb{R}^{C})$ parametrized by $\theta\in\Theta$, are trained on supervised ID data $\mathcal{D}^{n}\sim\bm{P}^{n}$ to output posteriors over logits corresponding to $C$ classes. For an input $x\in\mathcal{X}$ and its ground-truth label $y\in\mathcal{Y}$:
@@ 2.3 Algorithmic insights
D3M is free from training data. D3M is designed to monitor deployment-time input streams without requiring the training data of its base model. This is an important advantage shared by yu2019unsupervised , however not enjoyed by other baselines in Table 1, whose computation of $\phi$ statistics require maintaining agreement on training data. When the neural feature extractor $\operatorname{FE}$ is millions or billions of parameters, trained with a comparably large dataset, it becomes infeasible to store the training data within edge compute nodes in order to run the monitoring mechanism.
No tradeoff with prediction accuracy. Furthermore, D3M does not trade-off prediction accuracy as the mean prediction model from the PPD is not modified during the construction of $\Phi$ nor its deployment (Appendix C.1). Therefore, ID generalization theories remain applicable (vapnik1999overview, ; bartlett2002rademacher, ; zhang2016understanding, ; nagarajan2019deterministic, ; jiang2019fantastic, ; lee2020neural, ). In addition, recording a collection of ID maximum disagreement rates reinforces the robustness of the method. While single instances of maximum disagreement rate are subject to noise in the choice of the held-out validation set, the sampling of posterior logits $z_{i}^{(k)}$, and the number of drawn samples $K$, a collection $\Phi$ of maximum disagreement rates coupled with a quantile test softens the effect of noise by leveraging large sample statistics.
Less diverse samples compared to its fully Bayesian counterpart. Harrison et. al., 2024 harrison2024variational suggests that one may apply a variational Bayesian treatment to the last layer only, saving on computational costs of forward and backward propagating through a Bayesian neural network, and enjoy largely the same uncertainty estimation and OOD detection performance. However, the usage of VBLL in D3M is unorthodox with respect to the original work: by sampling posterior weights from a Bayesian model, we are hoping to land on a set of weights such that the resulting model strongly disagree with the base model. VBLL allows more efficient sampling compared to to a fully Bayesian network, though lacking diversity in the sampled predictions, as the variability comes from the last layer only. This results in $K$ sampled logits mostly agreeing with the mean prediction. To improve the diversity of sampling, we employ (1) temperature scaling and (2) the sampling of predictions from the categorical distribution for the computation of maximum disagreement rather than argmax labeling. We defer the study of this tradeoff between fully Bayesian and VBLL to Appendix C.5.
@@ 2.4 Implementing D3M
The composition $\operatorname{VBLL}_{\theta}\circ\operatorname{FE}_{\theta}:\mathcal{X}\to\mathcal{P}(\mathbb{R}^{C})$ is readily trained end-to-end on ID data of size $n$ by maximizing the ELBO. For rounds $t\in[T]$, on a large held-out ID validation set, we sample datasets $\mathcal{D}^{m}_{t}$ of size $m\ll n$ with replacement for calibration. $\mathcal{D}^{m}_{t}$ is processed in a single forward pass, producing $m$ independent $C$-dimensional Gaussian posteriors:
@@ 3.1 Experimental setup
Datasets. (1) The UCI Heart Disease prediction dataset (heart_disease_45, ), where each hospital corresponds to a different domain. Here, the distribution shift is due to differences in patient populations and data collection practices across hospitals. (2) CIFAR-10/10.1 datasets (krizhevsky2009learning, ; recht2019imagenet, ) where shift comes from subtle changes in the dataset creation process. By viewing samples from CIFAR-10 as $\bm{P}$, we test our models’ ability to flag deteriorating shift from samples in $\bm{Q}=$ CIFAR-10.1. (3) The Camelyon17 dataset from the WILDS benchmark (koh2021wilds, ; yao2022wild, ) a histopathology image dataset for detecting metastases in lymph node slides, where distribution shift arises from variations in slide staining and image acquisition between hospitals contributing the data. (4) The General Medicine Inpatient Initiative (GEMINI) (verma2017, ; verma2021assessing, ) dataset, a comprehensive repository of standardized clinical and administrative data from hospitalizations within general internal medicine. We focus on predicting patient mortality within a 14-day horizon, leveraging static and longitudinal clinical features. This task is deemed essential for facilitating timely clinical interventions, optimizing the allocation of healthcare resources, and ultimately striving to improve patient outcomes naemi2021machine ; barsasella2022machine ; hernandez2023machine . Detailed data descriptions can be found in B.3. Full sweeping details and hyperparameter configurations are reported in B.5.
Implementation & Baselines. For tabular data (UCI, GEMINI 14-day mortality), in our implementation, $\operatorname{FE}_{\theta}$ corresponds to sequences of affine transformations and nonlinearities with skip-connections. For image datasets (CIFAR-10/10.1, Camelyon17), $\operatorname{FE}_{\theta}$ is either a trained or a finetuned ResNet (he2016deep, ). In particular, the D3M mechanism is agnostic to the choice of feature extractor, provided that the architecture reflects appropriate inductive biases for the data modality and permits gradient flow through the extracted features.
To demonstrate that our D3M enjoys low FPR on non deteriorating shifts and high TPR on deteriorating shifts, we compare it against several distribution divergence-based detection methods from the literature: Deep Kernel MMD (MMD-D) (liu2020learning, ), H-divergence (zhao2022comparing, ), Black Box Shift Detection (BBSD) (lipton2018detecting, ), Relative Mahalanobis Distance (RMD) ren2021simple , Deep Ensembles (ovadia2019can, ), CTST (lopez2016revisiting, ), Domain Classifier (DC) (rabanser2019failing, ), and Detectron ginsberg2023a . Details can be found in Appendix B.1.
@@ 3.2 Discussions and Analysis
Comparisons with Detectron. ginsberg2023a While D3M and Detectron achieve similar monitoring performance—both successfully flagging all deteriorating shifts in higher query size scenarios (see B.6)—D3M is significantly more scalable and practical for real-world deployment, especially at the edge (Appendix C.2). Unlike Detectron, which requires persistent access to the original training data and gradient-based finetuning during deployment, D3M operates in a truly “source-free” fashion post-training: it needs neither storage nor access to sensitive training samples, nor does it perform computationally expensive or potentially destabilizing finetuning in production. Furthermore, by relying solely on forward passes and simple statistics over deployment data, D3M is inherently robust and less susceptible to the risks associated with continual retraining, such as accidental data contamination or privacy leakage.
@@ 4 Related Work
Adaptation. Test-time adaptation (TTA) concerns itself more with how one achieves strong deployment performance in spite of a critical change, rather than how one would flag this critical change. To this end, we fundamentally believe that monitoring offers an additional level of flexibility: when a deterioration flag is raised, adaptation is a possible course of action among many others. A few works from this rich body of literature are tangentially related to disagreement-based monitoring. saito2018maximum ; vu2019advent ; wang2020tent all leverage prediction uncertainty or discrepancy to drive unsupervised domain adaptation. zhang2023adaptive , measures and minimizes prediction disagreement at test time, adaptively aligning target-domain features to a domain-invariant feature space, improving performance without requiring access to target domain data during training. In the continual learning literature, ye2022task also leverages classifier disagreement as a unsupervised proxy of distributional change.
@@
Performance monitoring of ML models & deteriorating shift. Evaluating a model’s reliability during deployment is crucial for the safety and effectiveness of the machine learning pipeline over time. pmlr-v236-feng24a ; pmlr-v238-feng24a provided a causal viewpoint wherein the challenge to adapt to diverse scenarios still remain due to a lack of access to the true causal graph. Model disagreement is often used as a monitoring tool for the model generalization (jiangassessing, ; chuang2020estimating, ; ginsberg2023a, ; rosenfeld2024almost, ). These works align strongly with our method, though the framing is orthogonal to ours as our analysis provides FPR and TPR guarantees whereas they provided sufficient conditions in either ID or other shifts beyond our scope. Several works in the recent literature differentiate shifts in terms of deteriorating or non-deteriorating shifts. podkopaev2021tracking studied (deteriorating) shift detection in the continuous monitoring setting using a sequential hypothesis test. Due to the setting being sequential in nature, their method requires true labels from $\bm{Q}$ immediately after prediction or at the least in a delayed fashion. ginart2022mldemon approached the monitoring problem from a time-continuous anomaly detection perspective, allowing periodic querying of deployment-time ground truths from experts. Other related empirical works along this literature are Wang2023FeatureAE ; kamulete2022test , and the very applied nigenda2022amazon .
@@
Distribution shift detection. Methods to detect distribution shift arise from different perspectives. In covariate shift detection, (lopez2016revisiting, ; liu2020learning, ; zhao2022comparing, ) treated detection as two-sample tests via classifier, Deep Kernel MMD, and H-divergence. For label shift on the other hand, (lipton2018detecting, ; azizzadenesheli2019regularized, ) formulated the problem as a convex optimization problem by solving the label distribution ratio $\alpha=\bm{Q}(y)/\bm{P}(y)$. The problem of OOD detection (liang2017enhancing, ; kamulete2022test, ) seeks to detect if an individual sample $x$ comes from the training distribution $x\sim\bm{P}(x)$. Some previous works (ren2019likelihood, ; morningstar2021density, ) also adopted the methods in covariate shift detection and generalization by estimating the density ratio for the identification of OOD samples. Whilst these methods detect shifts, they are constrained by their requirement of training data post-deployment and do not consider the extent to which shifts affect model performance.
@@
Estimating test error with unlabeled data. Another rich body of research is the estimation of (OOD) test error. This technique and its variants are often inspired by domain adaptation theories (ben2006analysis, ; ben2010theory, ; acuna2021f, ; ganin2016domain, ), seeking guarantess in the form of $\operatorname{{err}}(f;\bm{Q}_{g})\leq\operatorname{{err}}(f;\bm{P}_{g})+\Delta(f,\mathcal{H})$, with $\Delta(f,\mathcal{H})=\sup_{h\in\mathcal{H}}|\operatorname{{err}}(h;\bm{P}_{f})-\operatorname{{err}}(h;\bm{Q}_{f})|$. This objective can be alternatively viewed as searching for a critic function $h\in\mathcal{H}$ to maximize a performance gap (rosenfeld2024almost, ; jiang2021assessing, ). One could thus provably estimate the upper bound of the test distribution error. These theories also implicitly assume the availability of training data. Further, they assume that test error should be larger than the training error (granted this is often the case for deteriorating OOD), making them sensitive to non deteriorating shifts as well i.e., a high FPR in detection. Our theoretical analysis addresses this gap.